Approximately 60% of all part-time college students in the Peshawar are female. What would

you expect to see in terms of the behavior of a sample proportion of females if random samples of

size 100 were taken from the population of all part-time college students? Estimate the standard

deviation also construct the confidence interval using alpha 0.05.


1
Expert's answer
2021-02-22T14:42:31-0500

As we know, due to sampling variability, sample proportion in random samples of size "100"

 will take numerical values which vary according to the laws of chance:

In other words, sample proportion is a random variable. To summarize the behavior of any random variable, we focus on three features of its distribution: the center, the spread, and the shape.


Based only on our intuition, we would expect the following:


Center: Some sample proportions will be on the low side — say, "0.55 \text{or} 0.58" — while others will be on the high side — say, "0.61 \text{or} 0.66" . It is reasonable to expect all the sample proportions in repeated random samples to average out to the underlying population proportion,

 "0.6" . In other words, the mean of the distribution of p-hat should be p.


Spread: For samples of "100" , we would expect sample proportions of females not to stray too far from the population proportion "0.6" .Sample proportions lower than "0.5" or higher than "0.7" would be rather surprising.

On the other hand, if we were only taking samples of size "10" , we would not be at all surprised by a sample proportion of females even as low as "\dfrac{4}{10} = 0.4," or as high as "\dfrac{8}{10} = 0.8." Thus, sample size plays a role in the spread of the distribution of sample proportion:there should be less spread for larger samples, more spread for smaller samples.


Shape: Sample proportions closest to "0.6" would be most common, and sample proportions far from "0.6" in either direction would be progressively less likely. In other words, the shape of the distribution of sample proportion should bulge in the middle and taper at the ends: it should be somewhat normal.


The standard deviation is given by the formula-


"\sigma_p=\sqrt{\dfrac{p(1-p)}{N}}"


"=\sqrt{\dfrac{0.6(1-0.6)}{100}}=\sqrt{\dfrac{0.24}{100}}" ="\dfrac{0.489}{10}=0.0489"



So The two confidence interval for "\alpha=0.05 \text{ is given by- }"


The probability is approximately 0.95 that p-hat falls within 2 standard deviations of the mean


=, "0.6-2(\alpha)=0.6 – 2(0.05) =0.5\text{ and } 0.6+2(\alpha)= 0.6 + 2(0.05)=0.7."


Hence, TheConfidence interval is (0.5,0.7)